### v1

# Annamaria Porreca Fabrizio Maturo Viviana Ventre

Abstract
Social network analysis (SNA) helps us understand the relationships and interactions between individuals, groups, organisations, or other social entities. In SNA, ties are generally binary or weighted based on their strength. Nonetheless, when actors are individuals, the relationships between actors are often imprecise and identifying them with simple scalars leads to information loss. Social relationships are often vague in real life. Despite many classical social network techniques contemplate the use of weighted links, these approaches do not align with the original philosophy of fuzzy logic, which instead aims to preserve the vagueness inherent in human language and real life. Dealing with imprecise ties and introducing fuzziness in the definition of relationships requires an extension of social network analysis to fuzzy numbers instead of crisp values. The mathematical formalisation for this generalisation needs to extend classical centrality indices and operations to fuzzy numbers. For this reason, this paper proposes a generalisation of the so-called Fuzzy Social Network Analysis (FSNA) to the context of imprecise relationships among actors. The article shows the theory and application of real data collected through a fascinating mouse tracking technique to study the fuzzy relationships in a collaboration network among the members of a University department.
keywords:
Fuzzy Social Network , Fuzzy Centrality Indices , Vague Relationships , Uncertainty , Fuzzy Numbers , University Department Collaboration Network.
MSC:
62A86 , 94D05 , 91D30
[…]
[inst1]organization=Department of Medical, Oral and Biotechnologies Science, University G. d’Annunzio of Chieti-Pescara, addressline=Via dei Vestini, 31, city=Chieti, postcode=66100, country=Italy
[…]
[inst2]organization=Faculty of Technological and Innovation Sciences, Universitas Mercatorum, addressline=Piazza Mattei, 10, city=Rome, postcode=00186, country=Italy
[…]
[inst3]organization=Department of Mathematics and Physics, University of Campania Luigi Vanvitelli, addressline=Viale Abramo Lincoln, 5, city=Caserta, postcode=81100, country=Italy
1 Introduction
Social network analysis (SNA) helps us understand the relationships and interactions between individuals, groups, organisations, or other social entities [18, 31, 29, 40, 26]. Each type of social structure can be represented with a graph, in which the social entities are represented as nodes or vertices in a network, and their relationships are represented as edges or ties. SNA examines relationships between entities and how information, resources, or influence flows through them. It identifies patterns, structures, and dynamics of social systems. Ties in a network are generally binary based on the presence or absence of the link or weighted based on their strength. Each connection can be identified by a number in $[0,1]$ or $\{0,1\}$, respectively. Nonetheless, when the actors of a social network are individuals, the relationships are very complex and measuring them using a single number leads to a loss of information. In other words, relationships are inherently imprecise in real life, and thus, social network analysis should consider vagueness. For example, let’s consider a social network like Facebook. The nodes may represent people registered on the network, and the links can be the connections between pairs of individuals. Typically, in a social network of this type, these connections are based on the presence or absence of ”friendship”. We must consider whether being friends with someone on Facebook should have the same significance for all connections. It’s evident that not all friendships are of equal weight and intensity; the level of interaction determines the closeness of a relationship rather than merely being connected online. However, this issue applies to social networks and real-life relationships because connections are fuzzy and real-life phenomena are not black or white but often grey [45, 47, 46]. What does it mean to be a friend? Is it being a close friend, an acquaintance, or just a casual acquaintance? What does being close collaborators or barely knowing each other in a university department mean?
One way to address this issue is using a fuzzy network analysis approach. Despite this, fuzzy network analysis has mainly been studied by introducing weighted ties within the network [see, e.g., 20, 8]. This approach of weighing the links between nodes differs from the original philosophy of fuzzy logic, which instead aims to preserve the vagueness inherent in human language. Assigning a link a value in $[0,1]$ or $\{0,1\}$ can sometimes be a simplification that sacrifices information for the sake of convenience. Translating vague concepts (such as close collaborator, occasional collaborator, real friend, and similar) necessarily results in a loss of information regarding the inherently imprecise nature of these expressions. Hence, considering a link as ”fuzzy” merely because it is represented by a number in the interval $[0,1]$ does not align with the basic principles of fuzzy logic, which begins and thrives where opposites coexist. In other words, fuzzy logic deals with situations where A and non-A coexist simultaneously with a certain degree of truth [45, 47, 46, 41]. Dealing with imprecise ties and introducing fuzziness in the definition of relationships needs an extension of classical social network analysis to fuzzy numbers instead of crisp values. The mathematical formalisation of this generalisation requires rewriting the classical centrality indices into fuzzy centrality indices involving fuzzy relationships. In other words, this research’s primary motivation and basic idea is that relationships between nodes should be characterised by functions rather than scalars to preserve the information about the vagueness of ties. The great advantage of this approach is the ability to create social networks and centrality indices that preserve information about the vagueness and imprecision of ties because they are inherent in human nature, human language, and interpersonal relationships. In statistics, it is known that valuable information is contained within variability. Similarly, in social networks, true information lies in understanding the vagueness of relationships.
The article shows the theory of Fuzzy Centrality Measures (FCM) in Social Network Analysis and their application to real data. The dataset adopted in this study is collected through a fascinating technique that includes mouse tracking to capture the vagueness of the relationships among the members of a University department. The first part of the work introduces the preliminaries necessary for formally understanding the proposed indices. Subsequently, based on the idea mentioned above, new centrality indices are proposed: fuzzy degree centrality, fuzzy out-degree centrality index, total fuzzy degree centrality index, fuzzy betweenness centrality, fuzzy closeness centrality, and fuzzy out-closeness centrality. The application Section proposes a fascinating investigation based on a University department collaboration network. The paper ends with a discussion and conclusions.
2 A brief introduction to Graph Theory (GT)
A graph $G=(V,E)$ is an abstract object formed by a set $V$ of vertices (nodes) and a set $E$ of edges (links) that join pairs of vertices. The vertex set and edge set of a graph $G$ are denoted by $V(G)$ and $E(G)$, respectively. The cardinality of $V$ is usually denoted by $n$, and the cardinality of $E$ by $m$. If an edge joins two vertices, they are adjacent, and we call them neighbors. Graphs can be undirected or directed. In undirected graphs, the order of the end vertices of an edge does not indicate a specific direction between them; the relationship is symmetric. In directed graphs, each directed edge (arc) has an origin (tail) and a destination (head). An edge with origin $u\in V$ and destination $v\in V$ is represented by an ordered pair $(u,v)$. In shorthand notation, an edge $\{u,v\}$ or $(u,v)$ can also be denoted by $uv$. In a directed graph, $uv$ is short for $(u,v)$, while in an undirected graph, $uv$ and $vu$ are the same and both stand for $\{u,v\}$. Edge weights are numerical values associated with edges or vertices in a graph $G=(V,E)$. Here, we focus on edge weights. These weights are often represented by a function $w:E\to\mathbb{R}$, assigning each edge $e\in E$ a weight $w(e)$. Depending on the application, edge weights can represent properties such as cost, capacity, strength of interaction, or similarity.
A walk from $v_{0}$ to $v_{k}$ in a graph $G=(V,E)$ is an alternating sequence $v_{0},e_{1},v_{1},e_{2},v_{2},$
$\ldots,v_{k-1},e_{k},v_{k}$ of vertices and edges, where $e_{i}=\{v_{i-1},v_{i}\}$ in the undirected case and $e_{i}=(v_{i-1},v_{i})$ in the directed case. The length of the walk is defined as the number of edges on the walk. The walk is called a path if $e_{i}\neq e_{j}$ for $i\neq j$, and a path is a simple path if $v_{i}\neq v_{j}$ for $i\neq j$. A path with $v_{0}=v_{k}$ is a cycle. A cycle is a simple cycle if $v_{i}\neq v_{j}$ for $0\leq i<j\leq(k-1)$. For a path $p$ in a graph $G=(V,E)$ with edge weights $w$, the weight of the path, denoted by $w(p)$, is defined as the sum of the weights of the edges on $p$. A path from $u$ to $v$ in $G$ is the shortest path (with respect to $w$) if its weight is the smallest possible among all paths from $u$ to $v$. The length of the shortest path from $u$ to $v$, also called the shortest-path distance between $u$ and $v$, is denoted by $d(u,v)$.
The single-source shortest paths problem is defined as follows: Given a graph $G=(V,E)$ with edge weights $w:E\to\mathbb{R}$ and a vertex $v\in V$ (the source), compute the shortest paths from $v$ to all other vertices in the graph. The problem is only well-defined if the graph does not contain a negative weight cycle. If the edge weights are non-negative, the shortest paths problem can be solved in time $O(m+n\log n)$ using an efficient implementation of Dijkstra’s algorithm [15]. If the edge weights are arbitrary, the Bellman-Ford algorithm [4] uses time $O(mn)$ to detect a cycle of negative length or, if no such cycle exists, solve the problem. For the case of unit edge weights, the Bellman-Ford algorithm solves the problem in linear time $O(n+m)$. In the all-pairs shortest paths problem, one is given a graph $G=(V,E)$ with edge weights $w:E\to\mathbb{R}$ and wants to compute the shortest-path distances for all pairs of nodes. Provided that $G$ does not contain a cycle of negative length, this problem can be solved by the Floyd-Warshall algorithm [17] in time $O(n^{3})$, or by running the shortest paths’ problem computations in time $O(nm+n^{2}\log n)$.
Adjacency and incidence matrices are tools used in graph theory to represent and analyze the relationships between vertices and arcs of a graph. The adjacency matrix (or sociomatrix) of a directed graph $G=(V,E)$ with vertices $V=\{v_{1},\ldots,v_{n}\}$ is an $n\times n$ matrix $A_{G}$ where $a_{i,j}=1$ if $(v_{i},v_{j})\in E$, and 0 otherwise. For undirected graphs, this matrix is symmetric. In weighted graphs, entries $a_{i,j}$ can represent weights $w(v_{i},v_{j})$ instead of binary adjacency. In a simple directed graph, the incidence matrix $B_{G}$ with $n$ rows and $m$ columns captures relationships between nodes and arcs. Entries $b_{i,j}$ indicate whether node $v_{i}$ is the origin or destination of arc $e_{j}$. A simple graph is a graph without loops (self-edges) and multiple arcs between any pair of vertices. For a simple direct graph, we can assume by convention $b_{i,i}=0$ $\forall i\in I_{n}$, for $i\neq j$.
3 Overview of the Main Classical Centrality Measures
When examining relationships within a group, our initial inquiry often revolves around identifying influential individuals within the network. To address such queries effectively, it is essential to establish measures that quantify centrality. Many vertex centrality indices were introduced in the 1950s, such as the Bavelas index [2, 3], degree centrality [21], and a first feedback centrality introduced by Seeley [36]. Before providing a formal definition of centrality indices, it is necessary to introduce the concept of isomorphism.
Definition 1
Two directed graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ are isomorphic (denoted as $G_{1}\simeq G_{2}$) if there exists a bijection $\varrho:V_{1}\rightarrow V_{2}$ such that for all $u,v\in V_{1}$:
$(u,v)\in E_{1}\Longleftrightarrow(\varrho(u),\varrho(v))\in E_{2}.$
Such a bijection is called an isomorphism.
An isomorphism that maps a graph onto itself is called an automorphism. Usually, we consider two graphs to be the same if they are isomorphic. Isomorphism and automorphism for a graph with symmetric relations are defined analogously.
Formally [24], a centrality index is a structural index which is defined as follows:
Definition 2 (Structural Index)
Let $G=(V,E)$ be a weighted, directed, or undirected multigraph, and let $V$ represent the set of vertices or edges of $G$, respectively. A real-valued function $s$ is called a structural index if and only if the following condition holds: for all $v\in V$, for any graph $H=(V^{\prime},E^{\prime})$ isomorphic to $G$, there exists a bijection $\varrho:V\rightarrow V^{\prime}$ such that $s_{G}(v)=s_{H}(\varrho(v))$, where $s_{G}(v)$ denotes the value of $s(v)$ in graph $G$.
A centrality index $C$ is required to be a structural index and thus induces at least a semi-order on the set of vertices or edges, respectively. According to this order, $v\in V$ is considered at least as central as $u\in V$ with respect to a given centrality $C$ if $C(v)\geq C(u)$. Centrality indices are structural measures used to quantify the importance of a node $v\in V_{G}$ in a graph $G$. For instance, $\sigma^{G}:V_{G}\rightarrow\Re_{0}^{+}$ assigns higher values to nodes that are more interconnected. In cases where two nodes $v$ and $u$ exhibit identical network structures, the centrality index should yield the same value $\sigma^{G}(v)=\sigma^{G}(u)$ for both nodes [28, 35]. Nodes $v$ and $u$ are considered identically structurally integrated into the network if there exists an automorphism $\eta:V_{G}\rightarrow V_{G}$ with $u=\eta(v)$. Centrality represents the extent to which an actor occupies a central position within a network. Empirical research has shown that occupying a central position in a network of relationships often correlates with access to resources and influence [14, 32, 33].
3.0.1 Degree
The number of edges connected to a vertex is referred to as its “degree” in graph theory. The degree of a node $v_{i}$ is commonly denoted as $d_{i}$. In the context of directed graphs, nodes have both in-degrees (edges pointing towards the node) and out-degrees (edges pointing away from the node), denoted as $d_{i}^{in}$ and $d_{i}^{out}$, respectively. For example, in social platforms like Facebook, the degree of a user corresponds to the number of friends [48].
Definition 3
In any undirected graph $G(V,E)$, the sum of all node degrees equals twice the number of edges:
$\sum_{i}d_{i}=2|E|$ (1)
This equation holds because each edge contributes to the degree sum of exactly two nodes.
In directed graphs, the sum of in-degrees equals the sum of out-degrees:
$\sum_{i}d_{i}^{in}=\sum_{j}d_{j}^{out}$ (2)
This equality arises because each directed edge contributes exactly one unit to the in-degree of its destination node and one unit to the out-degree of its origin node.
3.0.2 Degree centrality
In real-world interactions, individuals with many connections are often perceived as important. Degree centrality quantifies this idea into a measurable index. The degree centrality index ranks nodes higher based on the number of connections they have.
For an undirected graph, the degree centrality $C_{d}$ of a node $v_{i}$ is given by:
$C_{d}(v_{i})=d_{i}$ (3)
where $d_{i}$ represents the degree (number of adjacent edges) of node $v_{i}$.
In directed graphs, we distinguish between in-degree centrality, out-degree centrality, and total degree centrality, which combines in and out degrees:
[…]
In-degree centrality indicates how popular a node is within the network. Out-degree centrality measures how gregarious a node is.
Comparing degree centrality values across different networks directly is not straightforward due to variations in network size. To standardize comparison, degree centrality can be normalized as:
$C_{d}^{*}(v_{i})=\frac{d_{i}}{n-1}$ (7)
where $n$ is the number of vertices in the network.
3.0.3 Closeness centrality
Closeness centrality measures the average distance from a vertex to all other vertices in the network. Unlike other network metrics, closeness centrality offers a distinct perspective on each individual’s network position, capturing how quickly each vertex can reach all others [19]. The central idea is that nodes with higher closeness centrality are more central, allowing them to reach other nodes more efficiently [48]. Mathematically, closeness centrality $C_{c}(v_{i})$ for vertex $v_{i}$ is defined as:
$C_{c}(v_{i})=\frac{1}{\bar{l}_{D_{j}}}$ (8)
where
$\bar{l}_{D_{j}}=\frac{1}{n-1}\sum_{v_{j}\neq v_{i}}l_{i,j}$ (9)
represents the average shortest path length from node $v_{i}$ to all other nodes in the network. Nodes with smaller average shortest path lengths $l_{i,j}$ exhibit higher closeness centrality.
3.0.4 Betweenness centrality
Another perspective on centrality considers how crucial nodes are in connecting other nodes. For a node $v_{i}$, one approach is to quantify the number of shortest paths between other nodes that pass through $v_{i}$:
$C_{b}(v_{i})=\sum_{\begin{subarray}{c}s\neq v_{i}\neq t\end{subarray}}\frac{\sigma_{st}(v_{i})}{\sigma_{st}}$ (10)
Here, $\sigma_{st}$ denotes the total number of shortest paths from node $s$ to $t$ (also referred to as information pathways), and $\sigma_{st}(v_{i})$ represents the number of those paths that pass through $v_{i}$. Essentially, this metric evaluates how central $v_{i}$ is in facilitating connections between any pair of nodes $s$ and $t$, known as betweenness centrality.
To standardize betweenness centrality across networks, normalization is necessary. This involves calculating its maximum possible value:
$C_{b}^{*}(v_{i})=\frac{C_{b}(v_{i})}{\max_{s\neq t}C_{b}(v_{i})}$
The maximum value occurs when $v_{i}$ lies on all shortest paths from $s$ to $t$ for every pair $(s,t)$, i.e., $\forall(s,t),s\neq t\neq v_{i},\frac{\sigma_{st}(v_{i})}{\sigma_{st}}=1$.
4 A brief introduction to Fuzzy Numbers
The limit of bivalent logic is that it addresses the real problems as if they were white or black [25]. Understanding the critical point at which a thing becomes a non-thing undermines traditional logic, which is based on black and white, all or nothing, true or false; according to Aristotelian logic, an element belongs to a set or does not belong at all. Aristotelian thought is essentially based on the principles of identity, non-contradiction, and the excluded middle. Since ancient times, Mathematicians and Philosophers have tried to find a rational explanation for paradoxes to save the dichotomy assumption and avoid dealing with the grey that exists in all things. The first to use the term fuzzy to describe logic based on uncertainty and vagueness was Lotfi Zadeh. When Zadeh published his first article on “Fuzzy Sets”, he introduced the concept of “membership value” which, in contrast to the bivalent theory, considers the possibility of intermediate values of truth; in particular, the degree of membership can take any value in the range $[0,1]$[45]. Fuzzy logic contemplates the grey and starts where the contradictions inherent in things begin, where A and non-A melt to a certain extent. Thus, fuzzy logic essentially deals with uncertainty due to vagueness.
In classical set theory, a set comprises all the elements of the discourse universe that satisfy a given membership function. For a traditional set, which we define as crisp, to distinguish it from a fuzzy set, the membership function is Boolean; it associates to each element $x$ of the universe a value of “true” or “false” depending on whether or not $x$ belongs to the set. Formally, a crisp set has the following characteristic function:
$\mu_{A}(x)=\begin{cases}1&\text{if x $\in$ A}\\ 0&\text{if x $\notin$ A}\end{cases}$ (11)
The theory of fuzzy sets posits that an element can partially belong to a set, determined by a membership function that assigns real values in the interval $[0,1]$.
A fuzzy set A is defined by a set of ordered pairs and a binary relation as follows:
$\emph{A}=\{(x,\mu_{A}(x))\mid x\in A,\mu_{A}(x)\in[0,1]\}$ (12)
where $\mu_{A}(x)$ is a function, known as the membership function, that specifies the degree to which any element $x$ in $A$ belongs to the fuzzy set A. This implies a gradual transition from membership to non-membership [47].
Formally, a fuzzy set A is defined by its characteristic function:
$\mu_{A}:X\rightarrow[0,1]$ (13)
being $X$ the universe of definition.
4.1 Definition of a fuzzy number
A fuzzy number $\mu(x)$ is a special case of a fuzzy set; in fact it can be defined as a fuzzy set, defined on real number, with a normal and convex membership function, such that there exists at least one point where the membership function takes the value “one”; it is a very useful tool for working with imprecise numerical quantities [45]. A fuzzy number is a function having as domain the set of real numbers and with values in $[0,1]$:
$\mu:\mathbb{R}\rightarrow[0,1]$ (14)
such that the following characteristics apply:
  1. 1.
    Bounded support: there are two real numbers a and b, with $a\leq b$, called the endpoints of $\mu$, such that:
    $\begin{cases}\mu(x)=0&\text{for}\qquad x\not\in[a,b]\\ \mu(x)>0&\text{for}\qquad x\in(a,b);\end{cases}$ (15)
  2. 2.
    Normality: there are two real numbers $c$ and $d$, with $a\leq c\leq d\leq b$ such that:
    $\mu(x)=1\qquad\text{if and only if}\qquad x\in[c,d].$ (16)
  3. 3.
    Convexity: $\mu(x)$ is a function increasing in the interval $[a,c]$ and decreasing in the interval $[d,b]$;
  4. 4.
    Compactness: for every $\alpha\in(0,1)$, the set $x\in\mathbb{R}:\mu(x)=\alpha$} is a closed interval.
The set of the real numbers $x$ such that $\mu(x)>0$ is said the support of the fuzzy number, and the interval $[c,d]$ is the core or central part. The intervals $[a,c)$ and $(d,b]$ are, respectively, the left part and the right part. The real numbers ${\mu}_{L}=c-a$, ${\mu}_{C}=d-c$, and ${\mu}_{R}=b-d$ are the left, middle, and right spreads, respectively. Their sum ${\mu}_{T}=b-a$ is the total spread of the fuzzy number [23]. The choice of the type of fuzzy numbers is delicate as it influences the characteristics of the processes of fuzzification and defuzzification of the outputs and the inputs [47]. The fuzzification allows to calculate the degree of membership of each numerical value assumed by one input variable to each fuzzy set defined for it. The defuzzification, instead, starting from the result obtained in the fuzzy inference process, calculates a real value for the output variable. The main membership functions representing fuzzy variables are triangular (TFN), trapezoidal, and bell-shaped. In this work, we concentrate on TFNs. A triangular fuzzy number $A$ is defined by the following membership function:
$\mu_{A}(x)=\begin{cases}\frac{x-a_{l}}{a_{m}-a_{l}}&\text{for}\quad a_{l}\leq x\leq a_{m}\\ \frac{x-a_{r}}{a_{m}-a_{r}}&\text{for}\quad a_{m}\leq x\leq a_{r}\\ \ 0&\text{otherwise}.\end{cases}$ (17)
where $[a_{l},a_{r}]$ is the support and $a_{m}$ is the core. As illustrated in Fig.1 $a_{l}$ and $a_{r}$ are respectively the left and right endpoints, while $a_{m}$ is the point where the membership function is equal to one. However, a TFN is often indicated using a simpler notation as follows: $A=(a_{l},a_{m},a_{r})$.
Figure 1: Triangular fuzzy number.
4.2 Ranking fuzzy numbers
Given a universe set $X$, a fuzzy set $A_{i}$ on $X$ is defined as $\{(x,\mu_{A_{i}}(x))\mid x\in X\}$, where $\mu_{A_{i}}(x)$ denotes the membership degree of $x$ in $A_{i}$ [9]. The support of a fuzzy number is defined as:
$A_{i}=\operatorname{cl}\{x\in X\mid\mu_{A_{i}}(x)>0\},$
where $\operatorname{cl}$ represents the topological closure operator. Let $\mathcal{F}$ denote the set of all fuzzy numbers. Since a fuzzy number $A_{i}$ is fully characterized by its membership function values, denote this function by $A_{i}(x)$.
The primary ranking methods can be categorized into two types [9]. The first type maps fuzzy numbers directly onto the real line using transformation functions $M:\mathcal{F}\rightarrow\mathbb{R}$. These functions associate each fuzzy number with a real number and use the ordering $\geq$ on the real line. For most methods, a higher associated value indicates a higher rank:
$M(A_{i})\geq M(A_{j})\Rightarrow A_{i}\succeq_{M}A_{j},$
where $\succeq_{M}$ denotes the dominance relation induced by $M$.
The second type of ranking method generates fuzzy binary relations, defined as functions $M:\mathcal{F}\times\mathcal{F}\rightarrow[0,1]$, where $M(A_{i},A_{j})\in[0,1]$ measures the degree to which $A_{i}$ is greater than $A_{j}$. Accordingly, fuzzy numbers are ranked based on:
$M(A_{i},A_{j})\geq M(A_{j},A_{i})\Rightarrow A_{i}\succeq_{M}A_{j}.$
From the fuzzy binary relation among fuzzy numbers, certain procedures derive an ordering relation [38, 39]. One widely used ranking method is the Center of Gravity (CoG) of a fuzzy number, introduced by [30]:
$\operatorname{CoG}(A)=\frac{\int_{-\infty}^{\infty}xA(x)\,\mathrm{d}x}{\int_{-\infty}^{\infty}A(x)\,\mathrm{d}x}.$ (18)
There exist numerous other proposals in the literature for ranking fuzzy numbers [see, e.g., 42, 7, 16, 11, 12, 22, 1, 27]; however, ranking methods are secondary to our study focus and serve a functional role in our application. Hereafter, we employ the Center of Gravity, though our approach may extend to other strategies.
4.3 A brief introduction to Fuzzy Graphs (FGs)
Graphs are fundamental mathematical structures commonly used to model networks and relationships [5, 6]. Rosenfeld [34] introduced the concept of a fuzzy graph, applying fuzzy set theory to graphs. According to Rosenfeld [34], a fuzzy graph is defined as a graph composed of vertices and edges, where the edges have membership functions defined as follows:
Definition 4 (Rosenfeld [34])
A fuzzy graph $G=(\sigma,\mu)$ consists of functions $\sigma:S\rightarrow[0,1]$ and $\mu:S\times S\rightarrow[0,1]$, where for all $x,y\in S$, $\mu(x,y)\leq\min\{\sigma(x),\sigma(y)\}$.
This definition specifies that in a fuzzy graph, both vertices and edges have membership values in the interval $[0,1]$. Yeh and Bang [44] proposed an alternative formulation of fuzzy graphs tailored for clustering analysis:
Definition 5 (Yeh and Bang [44])
A fuzzy graph $G=(V,R)$ consists of a set of vertices $V$ and a fuzzy relation $R$ on $V$, where the edges between vertices have a membership function $\mu_{R}:V\times V\rightarrow[0,1]$.
Unlike Definition 4, Yeh and Bang’s formulation allows for both fuzzy vertices and fuzzy edges. They designed this version of fuzzy graphs to better suit clustering applications, providing a more flexible representation.
Since their inception, fuzzy graphs by Rosenfeld [34] and Yeh and Bang [44] have been extensively studied for various applications. Blue et al. [5, 6] categorized fuzzy graphs into different types, including:
  1. 1.
    Fuzzy sets of crisp graphs:
    $G=\{(G_{1},\mu(G_{1})),(G_{2},\mu(G_{2})),\ldots,(G_{n},\mu(G_{n}))\}$
    where each crisp graph $G_{i}$ is associated with a membership function $\mu(G_{i})$.
  2. 2.
    Crisp vertices and edges with fuzzy connectivity:
    $G=\{V,E\},\quad V=\{v_{1},v_{2},\ldots,v_{n}\},\quad E=\{(h_{i},\sigma_{i}),(t_{i},\tau_{i})\mid i=1,2,\ldots,m\}$
    The edges’ connectivity is fuzzy, where heads $h_{i}$ and tails $t_{i}$ have membership values in $[0,1]$, while vertices and edges themselves are crisp.
  3. 3.
    Fuzzy vertices and crisp edges:
    $G=\{V,E\},\quad V=\{(v_{1},\mu(v_{1})),(v_{2},\mu(v_{2})),\ldots,(v_{n},\mu(v_{n}))\},\quad E=\{e_{1},e_{2},\ldots,e_{n}\}$
    The graph $G$ consists of fuzzy vertices with membership functions and crisp edges.
  4. 4.
    Crisp graph with fuzzy weight:
    $G=\{V,E\},\quad V=\{v_{1},v_{2},\ldots,v_{n}\},\quad E=\{e_{1},e_{2},\ldots,e_{n}\}$
    where edges have fuzzy weights defined by:
    $w_{i}=\{(w_{i,1},\mu(w_{i,1})),(w_{i,2},\mu(w_{i,2})),\ldots\}$
These classifications reflect the varying degrees of fuzziness applied to graphs, ranging from fully crisp vertices and edges to combinations of fuzzy vertices or edges with crisp counterparts.
4.4 Fuzzy operations
A fuzzy operation of $n$ variables is a function,
$f:(v_{1},v_{2},\ldots,v_{n})\in[0,1]^{n}\rightarrow[0,1]$
The most commonly used binary fuzzy operations are the “fuzzy intersections” or “t-norms” and the “fuzzy union” or “t-conorms”.
4.4.1 The “fuzzy intersection” or “t-norms”
Definition 6
Let $f:[0,1]\times[0,1]\rightarrow[0,1]$, it is a “fuzzy intersection” or “t-norm” if the following properties hold:
  • 1.
    (Commutative) $f(a,b)=f(b,a)$;
  • 2.
    (Associative) $f(f(a,b),c)=f(a,f(b,c))$;
  • 3.
    (Monotonicity) $a\leq a^{\prime}\Rightarrow f(a,b)\leq f(a^{\prime},b);\quad b\leq b^{\prime}\Rightarrow f(a,b)\leq f(a,b^{\prime})$;
  • 4.
    (Unity as a neutral element) $f(a,1)=a;\quad f(1,b)=b$.
We then write $afb$ to indicate $f(a,b)$. Some examples of fuzzy intersections are:
  • 1.
    Standard intersection $\cap$: $a\cap b=\min(a,b)$;
  • 2.
    Algebraic product $\odot$: $a\odot b=ab$;
  • 3.
    Limited difference $\odot_{d}$: $a\odot_{d}b=\max(0,a+b-1)$;
  • 4.
[…]
The following theorems hold:
Theorem 1
For each $a,b\in[0,1]$, $a\Delta b\leq a\odot b\leq a\cap b$.
Theorem 2
If $\otimes$ is any fuzzy intersection $\implies$ $a\Delta b\leq a\odot b\leq a\cap b$.
Theorem 3
The standard intersection is the only fuzzy intersection $\times$ which satisfies the following property:
  • 1.
    (Idempotence) $\forall a\in[0,1],a\odot a=a$.
4.4.2 The “fuzzy union” or “t-conorms”
Definition 7
Let $f:[0,1]\times[0,1]\rightarrow[0,1]$, it is a “fuzzy union” or “t-conorm” if the following properties hold:
  • 1.
    (Commutative) $f(a,b)=f(b,a)$;
  • 2.
    (Associative) $f(f(a,b),c)=f(a,f(b,c))$;
  • 3.
    (Monotonicity) $a\leq a^{\prime}\Rightarrow f(a,b)\leq f(a^{\prime},b);\quad b\leq b^{\prime}\Rightarrow f(a,b)\leq f(a,b^{\prime})$;
  • 4.
    (Zero as a neutral element) $f(a,0)=a;\quad f(0,b)=b$.
We then write $afb$ to indicate $f(a,b)$. Some examples of fuzzy unions are:
  • 1.
    Standard union $\cup$: $a\cup b=\max(a,b)$;
  • 2.
    Algebraic sum ssubscriptdirect-$\oplus_{s}$: $a\oplus_{s}b=a+b-ab$;
  • 3.
    Limited sum $\oplus_{L}$: $a\oplus_{L}b=\min(1,a+b)$;
  • 4.
[…]
The following theorems hold:
Theorem 4
For each $a,b\in[0,1]$, $a\cup b\leq a\oplus_{s}b\leq a\oplus_{L}b\leq a\Gamma b$.
Theorem 5
If $\oplus$ is any fuzzy union $\implies$ $a\cup b\leq a\oplus b\leq a\Gamma b$.
Theorem 6
The standard union is the only fuzzy union $\oplus$ which satisfies the following property:
  • 1.
    (Idempotence) $\forall a\in[0,1],a\oplus a=a$.
Definition 8
A fuzzy intersection $\otimes$ and a fuzzy union $\oplus$ are said to be associated if it holds: $a\otimes b+a\oplus b=a+b$ (Grassmann relation).
Theorem 7
The following fuzzy intersection and fuzzy union pairs are associated with each other:
  • 1.
    Standard Intersection and Standard Union;
  • 2.
    Algebraic product and Algebraic sum;
  • 3.
    Limited Difference and Limited Sum;
  • 4.
    Drastic intersection and Drastic union.
4.4.3 Fuzzy n-air operations
Let $h:(v_{1},v_{2},\ldots,v_{n})\in[0,1]^{n}\rightarrow[0,1]$ be an $n$-ary fuzzy operation. In order for it to represent an average of the values $v_{1},v_{2},\ldots,v_{n}$, it is necessary to require some axioms:
  • 1.
    (M1) (Idempotence)
    $(v_{1}=v_{2}=\ldots=v_{n}=k\in[0,1])\rightarrow h(v_{1},v_{2},\ldots,v_{n})=k$;
  • 2.
    (M2) (Growth)
    $(u_{i}\leq v_{i},i=1,2,\ldots,n)\rightarrow h(u_{1},u_{2},\ldots,u_{n})\leq h(v_{1},v_{2},\ldots,v_{n})$;
  • 3.
    (M3) (Commutativity)
    If $(u_{1},u_{2},\ldots,u_{n})$ is a permutation of $(v_{1},v_{2},\ldots,v_{n})$ then $h(u_{1},u_{2},\ldots,u_{n})=h(v_{1},v_{2},\ldots,v_{n})$;
  • 4.
    (M4) (Continuity)
    $h$ is a continuous function on $[0,1]^{n}$.
Definition 9
Let $w=(w_{1},w_{2},\ldots,w_{n})$ be a vector formed by elements belonging to the interval $[0,1]$ and such that $w_{1}+w_{2}+\ldots+w_{n}=1$. For each tuple $(a_{1},a_{2},\ldots,a_{n})$ of elements of $[0,1]$, let $(b_{1},b_{2},\ldots,b_{n})$ be the permutation of $(a_{1},a_{2},\ldots,a_{n})$ satisfying the condition $b_{1}\geq b_{2}\geq\ldots\geq b_{n}$. The ordered weighted average of $a_{1},a_{2},\ldots,a_{n}$, weighted with the weight vector $w$ (OWA = Ordered Weighted Average), is the number:
$m_{w}(a_{1},a_{2},\ldots,a_{n})=b_{1}w_{1}+b_{2}w_{2}+\ldots+b_{n}w_{n}$
The following theorems hold:
Theorem 8
$m_{w}(a_{1},a_{2},\ldots,a_{n})\in[0,1]$.
Theorem 9
For any weight vector $w=(w_{1},w_{2},\ldots,w_{n})$ that is non-negative and such that $w_{1}+w_{2}+\ldots+w_{n}=1$, the operation $h:(a_{1},a_{2},\ldots,a_{n})\in[0,1]^{n}\rightarrow m_{w}(a_{1},a_{2},\ldots,a_{n})\in[0,1]$ satisfies properties (M1), (M2), (M3), and (M4).
Theorem 10
Let $w=(w_{1},w_{2},\ldots,w_{n})$ and $w^{\prime}=(w_{1},w_{2},\ldots,w_{i}+h,\ldots,w_{j}-h,\ldots,w_{n})$ with $h>0$, $i<j$, two non-negative weight vectors such that $w_{1}+w_{2}+\ldots+w_{n}=1$. Then for all $(a_{1},a_{2},\ldots,a_{n})\in[0,1]^{n}$, $m_{w}(a_{1},a_{2},\ldots,a_{n})\leq m_{w^{\prime}}(a_{1},a_{2},\ldots,a_{n})$.
Theorem 11
Let $w_{\text{max}}$ be the weight vector with the first component equal to 1 and all others zero, and let $w_{\text{min}}$ be the weight vector with the last component equal to 1 and all others zero. For any weight vector $w=(w_{1},w_{2},\ldots,w_{n})$ that is non-negative and such that $w_{1}+w_{2}+\ldots+w_{n}=1$, it holds that for all $(a_{1},a_{2},\ldots,a_{n})\in[0,1]^{n}$,
$m_{w_{\text{min}}}(a_{1},a_{2},\ldots,a_{n})\leq m_{w}(a_{1},a_{2},\ldots,a_{n})\leq m_{w_{\text{max}}}(a_{1},a_{2},\ldots,a_{n}).$
Definition 10
Let $\diamondsuit$ be a binary operation on $[0,1]$ which satisfies the commutative, associative, and monotonic properties. The $n$-ary operation generated by $\diamondsuit$ is the function:
$\diamondsuit_{n}:(v_{1},v_{2},\ldots,v_{n})\in[0,1]^{n}\rightarrow v_{1}\diamondsuit v_{2}\diamondsuit\ldots\diamondsuit v_{n}\in[0,1].$
We observe that based on the commutative and associative properties, $v_{1}\diamondsuit v_{2}\diamondsuit\ldots\diamondsuit v_{n}$ does not depend on the order in which the operations are carried out or on any permutations of the elements. Additionally, properties (M2) and (M3) apply. If $\diamondsuit$ is also idempotent (M1) and continuous (M4), then the $n$-ary operation generated by $\diamondsuit$ is a weighted ordered average (OWA). This occurs if there exists a weight vector $w=(w_{1},w_{2},\ldots,w_{n})$ such that:
$v_{1}\diamondsuit v_{2}\diamondsuit\ldots\diamondsuit v_{n}=m_{w}(v_{1},v_{2},\ldots,v_{n}).$
Theorem 12
If $\diamondsuit$ is the standard intersection $\cap$, then $\diamondsuit_{n}=v_{1}\cap v_{2}\cap\ldots\cap v_{n}$ is a weighted ordered average associated with the weight $w_{\text{min}}=(0,0,\ldots,1)$, i.e., $m_{w_{\text{min}}}(v_{1},v_{2},\ldots,v_{n})$.
Theorem 13
If $\diamondsuit$ is the standard union $\cup$, then $\diamondsuit_{n}=v_{1}\cup v_{2}\cup\ldots\cup v_{n}$ is a weighted ordered average associated with the weight $w_{\text{max}}=(1,0,\ldots,0)$, i.e., $m_{w_{\text{max}}}(v_{1},v_{2},\ldots,v_{n})$.
An example of a weighted ordered average not generated by a binary operation is the simple arithmetic average where $m_{w}(v_{1},v_{2},\ldots,v_{n})=\frac{v_{1}+v_{2}+\ldots+v_{n}}{n}$, and therefore $w=\left(\frac{1}{n},\frac{1}{n},\ldots,\frac{1}{n}\right)$.
In fact, if $\diamondsuit$ is the binary operation that associates the mean $\frac{a+b}{2}$ to each ordered pair $(a,b)$ of elements of $[0,1]$, then the commutative, monotonic, continuous, and idempotent properties hold, but the associative property does not.
4.5 Ordering and metric space structures of the OWA operators
4.5.1 Sorting structure
Let $W$ be the set of vectors $w=(w_{1},w_{2},\ldots,w_{n})$ formed by $n$ elements belonging to the interval $[0,1]$ such that $w_{1}+w_{2}+\ldots+w_{n}=1$. In $W$, an order $\leq$ can be defined for each pair $w=(w_{1},w_{2},\ldots,w_{n}),w^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},\ldots,w^{\prime}_{n})$ as follows:
$w\leq w^{\prime}\Leftrightarrow(\forall i\in\{1,2,\ldots,n-1\},w_{i}\leq w^{\prime}_{i})$ (19)
Theorem 14
The pair $(W,\leq)$ forms a lattice with at least $w_{\min}=(0,0,\ldots,1)$ and maximum $w_{\max}=(1,0,\ldots,0)$.
4.5.2 Structure of the metric space
Let $W$ be the set of vectors $w=(w_{1},w_{2},\ldots,w_{n})$ formed by $n$ elements belonging to the interval $[0,1]$ such that $w_{1}+w_{2}+\ldots+w_{n}=1$. Let $k=(k_{1},k_{2},\ldots,k_{n})$ be a vector satisfying $k_{1}>k_{2}>\ldots>k_{n}$, with $k_{1}=1$ and $k_{n}=0$. In $W$, a distance (or metric) $d$, associated with $k$, can be defined for each $w,w^{\prime}\in W$ as:
$d(w,w^{\prime})=k_{1}|w_{1}-w^{\prime}_{1}|+k_{2}|w_{2}-w^{\prime}_{2}|+\ldots+k_{n}|w_{n}-w^{\prime}_{n}|$ (20)
Some authors [43, 8] have proposed the expression
$k_{i}=\frac{n_{i}}{n-1}$ (21)
Thus, (20) can be rewritten as:
$d(w,w^{\prime})=\frac{\left[\sum_{i=1}^{n-1}(n-i)|w_{i}-w^{\prime}_{i}|\right]}{n-1}$ (22)
Theorem 15
Let $d$ be the distance defined by (20). The pair $(W,d)$ forms a metric space. Furthermore,
  • 1.
    (D1) $d(w,w^{\prime})\leq 1$;
  • 2.
    (D2) $d(w,w^{\prime})=1\iff\{w,w^{\prime}\}=\{w_{\min},w_{\max}\}$;
  • 3.
    (D3) For all $w\in W$, $d(w,w_{\min})+d(w,w_{\max})=1$.
According to Theorem 15, we can define a ”degree of proximity” $v(w,w^{*})$ of a vector $w$ from a fixed vector $w^{*}$ by setting $v(w,w^{*})=1-d(w,w^{*})$. In particular, we can consider:
  • 1.
    The proximity $v(w,w_{\max})$ of a vector $w=(w_{1},w_{2},\ldots,w_{n})$ from the vector $w_{\max}$, called orness (unionity). If we assume (21), it results in:
    $\text{orness}(w)=v(w,w_{\max})=\left[\sum_{i=1}^{n-1}\frac{(n-i)w_{i}}{n-1}\right]$ (23)
    With simple calculations, we find:
    $\text{orness}(w_{\max})=1,\quad\text{orness}(w_{\min})=0,\quad\text{orness}(w_{\text{med}})=0.5$
    where $w_{\text{med}}$ is the vector having all components equal to $1/n$.
  • 2.
    The proximity $v(w,w_{\min})$ of a vector $w=(w_{1},w_{2},\ldots,w_{n})$ from the vector $w_{\min}$, called andness (intersection). From (D3), it follows that:
    $\text{andness}(w)=1-\text{orness}(w)$ (24)
    If we assume (21), then it results in:
    $\text{andness}(w)=1-\left[\sum_{i=1}^{n-1}\frac{(n-i)w_{i}}{n-1}\right]$ (25)
  • 3.
    The mediality of a vector of weights $w$, defined as $v(w,w_{\text{med}})$, where $w_{\text{med}}$ is the vector having all components equal to $1/n$.
4.5.3 FOWA: The Fuzzy Ordered Weighted Averaging Operator
Recently, Chen and Chen [13] formulated the FOWA operator to deal with information expressed by fuzzy numbers.
Definition 11
[37] The fuzzy ordered weighted averaging operator, denoted by FOWA, is the mapping $\text{FOWA}:\widetilde{\mathscr{P}}(\mathbb{R}^{+})^{n}\rightarrow\widetilde{\mathscr{P}}(\mathbb{R}^{+})$, which has an associated weighting vector $(w_{1},\ldots,w_{n})$, such that
$\text{FOWA}(\tilde{x}_{1},\ldots,\tilde{x}_{n})=\bigoplus_{i\in[n]}(w_{i}\odot\tilde{x}_{(i)})$
where the operations $\oplus$ and $\odot$ are a t-conorm and a t-norm respectively, and ()($\cdot$)( ⋅ ) denotes a permutation on $[n]$ such that $\tilde{x}_{(1)}\leq\ldots\leq\tilde{x}_{(n)}$.
It is evident that the FOWA operator extends the OWA operator. Particularly, for a weight vector $w$, we have:
  • 1.
    If $\mathbf{w}=(1,0,\ldots,0)$, then FOWA $=\tilde{x}_{(1)}$ (fuzzy minimum).
  • 2.
    If $\mathbf{w}=(0,\ldots,0,1)$, then FOWA $=\tilde{x}_{(n)}$ (fuzzy maximum).
  • 3.
    If $w_{i}=\frac{1}{n}$ for all $i$, then FOWA $=$ FAA (fuzzy arithmetic averaging).
If FOWA concerns triangular fuzzy numbers, it is simply referred to as FTOWA.
5 Fuzzy Graphs: Generalization of Weighted Graphs
In this work, we propose the use of Fuzzy Graphs (FG) as a generalization of Weighted Graphs (WG). A WG is one in which edges are associated with weights. A WG can be represented as $G(V,E,W)$, where $W$ represents the weights associated with each edge ($|W|=|E|$). For an adjacency matrix representation, instead of 00 and $1$, we can use the weight associated with the edge. This saves space by combining $E$ and $W$ into one adjacency matrix $A$, assuming an edge exists between $v_{i}$ and $v_{j}$ if and only if $W_{i,j}\neq 0$.
In the literature, there are different definitions and proposals for the concept of fuzzy graph [see e.g. 49]. Here, we focus on graphs where only the edges are fuzzy. Thus, FGs are considered as a generalization of WG, where edges are associated with fuzzy numbers.
Definition 12
An undirected fuzzy social network is defined as a fuzzy relational structure $\widetilde{G}_{un}=(V,E,\widetilde{A}_{un})$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ is a non-empty set of actors or nodes, and
$\widetilde{A}_{un}=\left(\begin{array}[]{ccc}\widetilde{A_{11}}&\cdots&\widetilde{A_{1n}}\\ \vdots&\ddots&\vdots\\ \widetilde{A_{n1}}&\cdots&\widetilde{A_{nn}}\end{array}\right)$
is an undirected fuzzy relation on $V$.
Many fuzzy relations are directional. A fuzzy relation is directional if the ties are oriented from one actor to another. Thus, Hu et al. [20] defined the directed fuzzy social network (DFSN) as follows:
Definition 13
A directed fuzzy social network is defined as a fuzzy relational structure $\widetilde{G}_{dir}=(V,E,\widetilde{A}_{dir})$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ is a non-empty set of actors or nodes, and $\widetilde{A}_{dir}=\left(\begin{array}[]{ccc}\widetilde{A_{11}}&\cdots&\widetilde{A_{1n}}\\ \vdots&\ddots&\vdots\\ \widetilde{A_{n1}}&\cdots&\widetilde{A_{nn}}\end{array}\right)$ is an undirected fuzzy relation on $V$.
Fuzzy Social Network (FSN) includes Undirected Fuzzy Social Network (UFSN) and Directed Fuzzy Social Network (DFSN). The significant difference between DFSN and UFSN is that a directed fuzzy relation is considered. According to Definition 12, $\widetilde{A}_{ij}$ is equal to $\widetilde{A_{ji}}$ in an undirected fuzzy social network. However, $\widetilde{A}_{ij}$ is not always equal to $\widetilde{A}_{ji}$ in a directed fuzzy social network. In Definition 13, $\widetilde{A}_{dir}$ is called the directed fuzzy adjacency matrix of $\widetilde{G}_{dir}$:
$\mu_{\widetilde{A}_{dir}}=\left(\begin{array}[]{ccc}\mu\left(\widetilde{A_{11}}\right)&\cdots&\mu\left(\widetilde{A_{1n}}\right)\\ \vdots&\ddots&\vdots\\ \mu\left(\widetilde{A_{n1}}\right)&\cdots&\mu\left(\widetilde{A_{nn}}\right)\end{array}\right)$
where, in this case, we have an asymmetrical membership function. The related concepts of UFSNs can also be used in DFSNs, but the direction of the fuzzy relation between actors must be considered.
Definition 14
Assume that $v_{0}A_{1}v_{1}A_{2}v_{2}\ldots A_{k}v_{k}$ is a path from $v_{0}$ to $v_{k}$ in $\widetilde{G}_{dir}$, then $\widetilde{s}_{d}(\widetilde{\omega})=\bigwedge_{i=1}^{k}\mu\left(A_{i}\right)$ is called the directed fuzzy intensity of path $\widetilde{\omega}$.
In Definition 2.3, $\mu(e_{i})$ is a membership function.
Definition 15
If $\bar{\omega}_{k}$ are $n$ paths from $u$ to $v$ in $\widetilde{G}_{dir}$, then $\widetilde{s}_{d}(u,v)=\bigvee_{k=1}^{n}\widetilde{s}_{d}(\widetilde{\omega}_{k})$ is called the directed fuzzy connected intensity from $u$ to $v$.
Here, $k=1,2,\ldots,n$. If there is no path from $u$ to $v$, then $\widetilde{s}_{d}(u,v)=0$. If $u=v$, then $\widetilde{s}_{d}(u,v)=1$. In UFSN, $\bar{s}(u,v)$ always equals $\widetilde{s}(v,u)$. However, $\widetilde{s}_{d}(u,v)$ is not always equal to $\widetilde{s}_{d}(v,u)$ in DFSN.
Definition 16
Assume $G_{dn}=(V,E_{dn})$ is a DFSN,
$\widetilde{C}_{d}=\left(\begin{array}[]{ccc}\widetilde{s}\left(v_{1},v_{1}\right)&\cdots&\widetilde{s}\left(v_{1},v_{n}\right)\\ \vdots&\ddots&\vdots\\ \widetilde{s}\left(v_{n},v_{1}\right)&\cdots&\widetilde{s}\left(v_{n},v_{n}\right)\end{array}\right)$
then $\widetilde{C}_{d}$ is called the directed fuzzy connected intensity matrix of $\widetilde{G}_{dn}$.
The directed fuzzy connected intensity matrix is a crucial concept in DFSN. With it, we can determine the directed fuzzy connected intensity between any two actors in DFSN, indicating the relationship (direct or indirect) between them.
Hereafter, we will focus only on triangular fuzzy numbers using a simple notation like the following:
$\tilde{A}=(a_{l},a_{m},a_{r}).$
An example of FG can be done to represent relations in a social network like Facebook or in a University Department, where the nodes are the department members, and the edges are the relations between them assessed according to a questionnaire. If we collect information about the relationship between pairs of Professors and researchers according to how much they feel they have relations with another member (for example, not at all, a little, a lot, etc.), summarizing this information with a scalar, as we do in the WG, would lead to a loss of information. Indeed, when dealing with this type of opinion, mainly when it is turned off through a human language attribute, the data always has a degree of inaccuracy. In such a context, it is natural to consider that the FG is direct because the opinion of member $A$ about his relationship with member $B$ may differ from that of member B about his relationship with member $A$. For this reason, we consider only the case of direct graphs. We will have the symmetrical FG be only a particular case of the asymmetrical one if all couples give the same opinion. This latter case would be improbable since we deal with personal opinions.
Assuming that through a fuzzy inference procedure, we have translated the qualitative answers of the department members into fuzzy numbers, we could have a situation like the one illustrated in Figure 2.
Figure 2: Example of fuzzy social network representing relationship in a University Department. The edges are represented via triangular fuzzy numbers. For simplicity, we do not plot the fuzzy number when there is a null relationship.
6 Fuzzy Centrality Indexes in a Fuzzy Social Network
Centrality indices quantify the importance of a node in a given network, which is often identified with the importance of the corresponding entity in the complex system modelled by the network. In the following sections, we extend the classical centrality indexes to the case of a network where the edges are represented by triangular fuzzy numbers. In other words, we define fuzzy degree centrality, fuzzy betweenness centrality, fuzzy closeness centrality, and fuzzy eigenvector centrality.
6.1 Fuzzy degree centrality
In real-world interactions, we often consider people with many connections to be important. Degree centrality transfers the same idea into a measure. The degree centrality measure ranks nodes with more connections higher in terms of centrality. In the classical framework, the degree centrality $C_{d}(v_{i})$, in directed graphs can be computed in different ways according to the number of in-connections and out-connections, leading to the in-degree index $d_{i}^{in}$ and the out-degree index $d_{i}^{out}$, respectively, or the combination of both. However, in the context of (directed) fuzzy social networks, the edges are represented by fuzzy numbers, and thus they are not scalars. Therefore, in the next Sections, we extend the concepts of in-degree centrality and out-degree centrality to the case of fuzzy numbers.
The first step is introducing the concepts of fuzzy triangular ordered weighted arithmetic operator [37]. The Fuzzy Triangular Ordered Weighted Arithmetic Operator (FTOWA) can be defined as follows.
Definition 17
The fuzzy triangular ordered weighted arithmetic operator, denoted by FTOWA, is the mapping FTOWA: $\widetilde{\mathscr{P}}([0,1])^{n}\rightarrow\widetilde{\mathscr{P}}([0,1])$, such that
$\operatorname{FTOWA}\left(\tilde{x}_{1},\ldots,\tilde{x}_{n}\right)=\bigoplus_{i\in[n]}\left(w_{i}\odot\tilde{x}_{(i)}\right)$
where $\oplus$ is a t-conorm, $\odot$ is a t-norm, $\left(w_{1},\ldots,w_{n}\right)$ is a weighting vector associated with $\oplus$, and ()($\cdot$)( ⋅ ) is a permutation on $[n]$ such that $\tilde{x}_{(1)}\leq\cdots\leq\tilde{x}_{(n)}$
6.1.1 Fuzzy in-degree centrality index
When using in-degrees, fuzzy degree centrality measures how popular a node is and its value shows prominence or prestige.
Definition 18
Assume that $\widetilde{G}_{dir}=\left(V,E,\widetilde{A}_{dir}\right)$ is a DFSN, $\widetilde{D}_{in}\left(v_{i}\right)$ is the sum of the fuzzy relations that are adjacent to $v_{i}$, then $\widetilde{D}_{in}\left(v_{i}\right)$ is called fuzzy in-degree centrality of $v_{i}$.
For a given node $i$, let $\widetilde{O}_{j}^{in}$ be the set of fuzzy numbers $\widetilde{A}_{ji}$ ranked in decreasing order, indicating the relationship between each node $j$ and the node $i$ (directed), the formula of fuzzy in-degree centrality of $v_{i}$ is given by
$\widetilde{D}_{in}\left(v_{i}\right)=\bigoplus_{j\in[n],j\neq i}\left(w_{j}\odot\tilde{O}_{j}^{in}\right).$ (26)
To rank the fuzzy numbers, we can use one of the approaches presented in Section 2.12. In the following, we will focus on the centre of gravity approach to understand the ranking of fuzzy numbers.
It follows that we can obtain the following extreme cases:
  • 1.
    $w_{j}=\{1,0,...,0\}$ $\implies\widetilde{D}_{in}$ is determined by the greatest $\widetilde{A}_{ji}$;
  • 2.
    $w_{j}=\{0,0,...,1\}$ $\implies\widetilde{D}_{in}$ is determined by the lowest $\widetilde{A}_{ji}$;
  • 3.
    $w_{j}=\{1/n,1/n,...,1/n\}$ $\implies\widetilde{D}_{in}$ is determined by all the $\widetilde{A}_{ji}$ with the same weights. Therefore, we give the same importance to low and high fuzzy relationships in computing fuzzy in-degree centrality.
The index just illustrated in Equation 26 does not allow us to compare the centrality values between different networks. To overcome this drawback, we can normalize the index so that it becomes a pure number between 0 and 1. The reasoning behind the normalization is that we calculate the maximum that the in-degree centrality index can take. The maximum of this index would, of course, be obtained if an actor is perfectly connected with all the other nodes. For this aim, the edge represented by the fuzzy number should have the characteristic of possessing a core with the maximum possible value on the support of the fuzzy number (with zero fuzziness). If the latter happens for all the actor’s edges, then we can compute the maximum of Equation 26 as follows.
The easiest way to find this relative index is to normalize the fuzzy numbers. In fact, unlike the classical case in which the number of edges is necessary to normalize the in-degree centrality, in this case, all the nodes are connected and therefore, what matters is the intensity of the edges. Once the fuzzy numbers have been normalized, we already have a relative index as follows:
$\widetilde{D}^{in*}\left(v_{i}\right)=\bigoplus_{j\in[n],j\neq i}\left(w_{j}\odot\widetilde{O}_{j}^{in*}\right).$ (27)
where $\widetilde{O}_{j}^{in*}$ is the order set of fuzzy numbers with normalized support in $[0,1]$.
6.1.2 Fuzzy out-degree centrality index
When using out-degrees, it measures the gregariousness of a node.
Definition 19
Assume that $\widetilde{G}_{dir}=\left(V,E,\widetilde{A}_{dir}\right)$ is a DFSN, $\widetilde{D}^{out}\left(v_{i}\right)$ is the sum of the fuzzy relations that are adjacent from $v_{i}$, then $\widetilde{O}_{j}^{out}\left(v_{i}\right)$ is called fuzzy out-degree centrality of $v_{i}$
For a given node $i$, let $\widetilde{O}_{j}^{out}$ be the set of fuzzy numbers $\widetilde{A}_{ij}$ ranked in decreasing order, indicating the relationship between each node $i$ and the node $j$ (directed), the formula of fuzzy out-degree centrality of $v_{i}$ is given by
$\widetilde{D}^{out}\left(v_{i}\right)=\bigoplus_{j\in[n],j\neq i}\left(w_{j}\odot\tilde{O}_{j}^{out}\right).$ (28)
According to the vector $w_{j}$, we can obtain the following cases:
  • 1.
    $w_{j}=\{1,0,...,0\}$ $\implies\widetilde{D}_{out}$ is determined by the greatest $\tilde{A}_{ij}$;
  • 2.
    $w_{j}=\{0,0,...,1\}$ $\implies\widetilde{D}_{out}$ is determined by the lowest $\tilde{A}_{ij}$;
  • 3.
    $w_{j}=\{1/n,1/n,...,1/n\}$ $\implies\widetilde{D}_{out}$ is determined by all the $\tilde{A}_{ij}$ with the same weights. Therefore, we give the same importance to low and high fuzzy relationships in computing fuzzy out-degree centrality.
Naturally, a possible solution can also be to adopt a vector of decreasing weights so that the importance of fuzzy numbers decreases as the relationship is less important.
As for the in-degree centrality index, also in this case, we can define a relative out-degree centrality index as follows:
$\widetilde{D}^{out*}\left(v_{i}\right)=\bigoplus_{j\in[n],j\neq i}\left(w_{j}\odot\tilde{O}_{j}^{out*}\right).$ (29)
6.1.3 Total Fuzzy degree centrality index
In DFSN applications, the previous degrees can be of great interest. The fuzzy out-degree centralities are measures of expansiveness and the fuzzy in-degree centralities are measures of receptivity, or popularity.
However, we can consider a formula for the total degree centrality of $v_{i}$ given by:
$\widetilde{D}\left(v_{i}\right)=\widetilde{D}^{out}\left(v_{i}\right)\oplus\widetilde{D}^{in}\left(v_{i}\right)$ (30)
6.2 Fuzzy betweenness centrality
Another way of looking at centrality is by considering how important nodes are in connecting other nodes. In the classical setting, for a node $v_{i}$, we can compute the relative frequency of the shortest paths to connect other nodes that pass through $v_{i}$.
In the classic case, the search for the shortest path consists of finding the path from one node to another node that is as fast as possible, crossing the least number of intermediate nodes. In our case, however, the nodes of the social network are all connected to each other with a different degree of truth; consequently, the shortest path, in a case where all the fuzzy edges would be equal, would always be the directed path from one node to another node. But usually, in reality, the connections are not all the same.
At this point in our setting, we are forced to reinterpret the concept of the shortest path as follows: In a context of fuzzy social network analysis, the importance of a node in connecting the other nodes must be evaluated based on the quality of the existing fuzzy edges: one path must be considered privileged over another if it is more effective in connecting two elements.
In our context, a path can be considered more effective than another path if it is able to connect two nodes, ensuring a higher connection strength. Clearly, it may happen that a longer path in terms of intermediate nodes to cross is better in terms of the strength of the edges and, therefore, more effective in connecting two elements $i$ to $j$. Therefore, in a network where everyone is (“more” or “less”) connected, an indirect path to go from node $i$ to $j$ could be better than the directed path from $i$ to $j$ (for example $i$ and $j$ know each other little directly but are well-connected through another subject).
Definition 20
Given two different paths $P_{i\rightarrow j}^{(1)}$ and $P_{i\rightarrow j}^{(2)}$, of possible different lengths $k={1,2,...,K}$ and $t={1,2,...T}$, respectively, we say that $P_{i\rightarrow j}^{(1)}$ is better than $P_{i\rightarrow j}^{(2)}$ when:
$P_{i\rightarrow j}^{(1)}\succ P_{i\rightarrow j}^{(2)}\iff\bigoplus_{k\in[1,K]}\left(w_{k}\odot\tilde{N}_{k}^{(1)}\right)>\bigoplus_{t\in[1,T]}\left(w_{t}\odot\tilde{N}_{t}^{(2)}\right)$ (31)
where $\widetilde{N}_{k}^{(1)}$ is the set of $K$ fuzzy edges in the first path between $i$ and $j$ with normalized support in $[0,1]$ and ranked in decreasing order, $\widetilde{N}_{t}^{(2)}$ is the set of fuzzy numbers in the second path between $i$ and $j$ also with normalized support in $[0,1]$ and ranked in decreasing order, and $w_{k}$ and $w_{t}$ is a vector of weights. We define $P_{i\rightarrow j}^{(1)}$ as the best path from $i$ to $j$ when for any other path $P_{i\rightarrow j}^{(2)}$ from $i$ to $j$ we have that $P_{i\rightarrow j}^{(1)}\succ P_{i\rightarrow j}^{(2)}$. Equation 31 allows us to consider different possibilities for selecting the so-called best path, e.g. average fuzzy number, minimum, maximum, etc. From now on, we will use the minimum by introducing $w_{k}=\{0,0,...,1\}$ and $w_{t}=\{0,0,...,1\}$ as the vectors of weights. In other words, when looking at two possible alternative paths to reach $j$ starting from the node $i$, we examine all the fuzzy edges in those paths, and we select the best one with the higher minimum fuzzy number (edge) in the path.
An example of this reasoning is illustrated in Figure 3. The best path for going from Dr. Perry to Dr. Porreca is via Prof. Smith. In fact, if we consider the direct connection, we have a very weak relationship, while if we connect the two nodes indirectly, we have two very strong relationships. In other words, if the direct relationship is weak, it is better to take more steps characterized by strong relationships. In the case of the example shown, the same applies if Dr. Porreca wants to get in touch with Perry (reverse path). This graphic example helps to better understand that in our context where the links are blurred, the concept of the shortest path is replaced by the concept of the best and most effective path to connect two distinct nodes. So we can say that the concept of betweenness centrality we have in the classic social network, that is, that an actor is highly central when he is in the middle of many shorter paths to go from a generic $j$ node to a generic $i$ node, in this context, it is replaced by the following concept: a node $z$ has high centrality if, considering all the best paths to go from a generic node $j$ to a generic node $i$, $z$ is frequently in the middle of these paths.
Figure 3: Example of fuzzy social network where the directed path between two nodes is worse than an indirected path in terms of fuzzy numbers edges. In other words, if the direct relationship is weak it is better to take more steps characterized by strong relationships.
Having identified the concept of the best path, we can therefore formalize the betweenness centrality of node $i$ as the ratio between all the best paths to join two nodes $s$ and $t$ via $i$ (obviously they are indirect paths), and all the best paths between all the pairs of nodes $s$ and $t$ (direct or indirect paths via $i$ or not). Formally, we can write as follows:
$B\left(v_{i}\right)=\sum_{s\neq t\neq v_{i}}\frac{P_{st}(v_{i})}{P_{st}}$ (32)
where $P_{st}(v_{i})$ indicate the number of “best” paths between $s$ and $t$ that passes thought $i$, and $P_{st}$ the total number of “best” paths between generic $s$ and $t$.
The search of the so-called “best” paths is done according to the minimum fuzzy number edge occurring in the path between $s$ and $t$, even if it is indirect. Clearly, to compare the multiple different paths that occur in a social network, we have to rank them according to the mentioned criteria to order fuzzy numbers, e.g. the center of gravity approach.
6.3 Fuzzy closeness centrality
In classical closeness centrality, the intuition is that the more central nodes are, the more quickly they can reach other nodes. Formally, these nodes should have a smaller average shortest path length to other nodes. The smaller the average shortest path length, the higher the centrality for the node.
As we have seen for fuzzy betweenness centrality, also in this case we refer to the concept of “shortest” path as “best” path, i.e. to reach $j$ starting from $i$, one path could better than another one despite the number of intermediate nodes is higher when the relationship between them is strong.
6.3.1 Fuzzy in-closeness centrality
Let the average (fuzzy number) of the best paths from node $i$ to all the nodes $j$ be represented by:
$\widetilde{M}_{v_{i},s}=\bigoplus_{v_{j}\neq v_{i}\in V}\left(w_{j}\odot\widetilde{A}_{ij}\right)$ (33)
where $\widetilde{A}_{ij}$ is the set of the best fuzzy edges to connect $i$ and $j$ ranked in decreasing order, s indicates the maximum number of steps considered evaluating the best path, and $w_{j}$ a vector of weights of the FOWA operator. Clearly, being a weighted average of fuzzy numbers, $\widetilde{M}_{v_{i},s}$ is also a fuzzy number.
Hence, we can define the fuzzy out-closeness centrality index as follows:
$\widetilde{C}^{in}_{v_{i},s}=\widetilde{M}_{v_{i},s}$ (34)
Thus, differently from Equation 32 that is a relative frequency, $\widetilde{C}_{v_{i},s}$ is a centrality index that incorporates the fuzziness of the edges and thus remains fuzzy.
Another possible approach to define fuzzy in-closeness centrality, in a way that is more close to the classical index, would be to consider:
$\widetilde{C^{(2)}}^{in}_{v_{i},s}=\frac{1}{\widetilde{M}_{v_{i},s}}.$ (35)
However, in the following we refer to Equation 34 because our basic idea is to look for an immediate representation of the in-closeness concept rather than to identically reproduce the classical indexes using fuzzy numbers. In fact, Equation 34 is quite interpretable, because provides a measure of how a node is central in considering the fuzzy incoming edges. However, Equation 34, differently from 26 takes into account the concept of the best path and maximum number of steps to reach another node.
6.3.2 Fuzzy out-closeness centrality
Following the same reasoning, we can define the fuzzy in-closeness centrality index of the node $v_{i}$ as follows:
$\widetilde{C}^{out}_{v_{i},s}=\widetilde{Z}_{v_{i},s}$ (36)
where $\widetilde{Z}_{v_{i},s}$ is the average (fuzzy number) of the best paths from nodes $j$ to $j$ in decreasing order. Thus,
$\widetilde{Z}_{v_{i},s}=\bigoplus_{v_{j}\neq v_{i}\in V}\left(w_{j}\odot\widetilde{A}_{ji}\right)$ (37)
is computed by applying the FOWA operator to fuzzy numbers representing the edges from the generic nodes $j$ and the node $i$, which we consider centrality.
As before, the vector of weights can be of different types. From now on, we will compute the average fuzzy number $\widetilde{M}_{v_{i},s}$ or $\widetilde{Z}_{v_{i},s}$ using $w=\{1/n,...,1/n\}$. However, we stress that to select the best paths $P_{i\rightarrow j}$ or $P_{j\rightarrow i}$ will always consider $w=\{0,...,1\}$. Naturally, we could also consider different combinations of the weights according to the different meanings we aim to give to our application. In other words, we apply the FOWA operator twice when computing betweenness centrality indexes.
As for the fuzzy in-closeness centrality index, also in this case we can consider an alternative measure of the fuzzy out-closeness centrality index as follows:
$\widetilde{C^{(2)}}^{out}_{v_{i},s}=\frac{1}{\widetilde{Z}_{v_{i},s}}$ (38)
In this case, we get an index more similar to the classical non-fuzzy in-closeness centrality index. The interpretation is contrary to 36; the lower the index, the higher the fuzzy in-closeness centrality index. Therefore, in the following, we focus on Equation 36 because the interpretation is quite immediate; indeed, the higher the index in Equation 36, the higher the strength of the fuzzy outcoming edges of a node (given the number of steps to consider the best path), the higher the fuzzy out-closeness centrality of that node.
6.3.3 Total fuzzy closeness centrality
As we have seen for fuzzy degree centrality, we can also sum up the two different components of closeness centrality in this case. Therefore, we can define the total fuzzy closeness centrality index as follows:
$\widetilde{C}^{tot}_{v_{i},s}=_{s}\tilde{C}^{in}_{v_{i},s}+_{s}\tilde{C}^{out}_{v_{i},s}$ (39)
7 An application of fuzzy social network analysis to collaboration networks in University Departments
Collaboration is a complex and multifaceted concept that cannot be evaluated using a single metric. Attempting to measure it directly takes a lot of work. One possible approach is to use proxies, such as counting the number of shared papers or similar information, but this method still needs to fully capture the essence of collaboration. Scientific collaboration is a broad concept involving exchanging knowledge and ideas between researchers. It includes sharing research procedures and ideas, which ultimately lead to the production of scientific knowledge. However, this definition alone is not enough to help us empirically measure a collaboration network of this type.
For this reason, this study starts from the basic idea of collecting data using the intuition of previous research in a different context. Notably, Calcagnì and Lombardi [10] proposed a new method to collect qualitative data using fuzzy set theory: the so-called Dynamic Fuzzy Rating Tracker (DYFRAT). The authors developed an applet to collect directly triangular fuzzy numbers data without the fuzzy inference process [10]. Their work used mouse movements to model human rating evaluations from a fuzzy-set perspective, capturing the imprecision in point answers. In particular, DYFRAT captures the fuzziness of human ratings by modelling some real-time biometric events that occur during the cognitive rating process the noisy and dynamic x-y trajectory is initially mapped into a set of polar objects (distances and angles). Next, a linear histogram model of angles of movements is built, which stores the most important spatial features of the computer mouse trajectory, such as location, directions, and amplitudes. Finally, the histogram model is used to determine a quantification of spatial events involved in the original movement path. DYFRACT is not available, and thus, in this paper, following the intuition by [10], we built an interactive fuzzy questionnaire to collect fuzzy numbers data to quantify and measure the collaboration relations between professors in the Economic Studies Department of the University “G. d’Annunzio” in Italy.
The software implemented to collect fuzzy data is based on the LAMP platform, i.e., Linux (operating system) + Apache (webserver) + MySql (database management system) + PHP (server-side programming language). For the management and use of the graphical interface was used Bootstrap (free framework initially created by Twitter to standardize graphical components, based in turn on HTML + CSS + JS) with its many modules, while also retrieving the position of the mouse, were used: JS (JavaScript, object-oriented programming language and client-side events) + jQuery (development library for client-side web applications based on JS).
Figure 4 shows an example of a single questionnaire answer inserted on a semicircumference. Each user is called to answer the same question with respect to her/his collaboration relay with all the other members of the department. This pseudo-circular scale is justified by its ability to provide greater degrees of freedom for mouse movement recordings than traditional linear or arc-type scales. All the information related to the questions is saved directly on Database. At the same time, the users’ answers are stored on DB in JSON format (JS-based information interchange language).
Figure 4: Fuzzy questionnaire to collect mouse movements of respondents.
The distribution of mouse movements within the semicircumference, through the tracking procedure, allowed us to grasp the uncertainty in the answer before clicking on the final answer. This allowed us to build triangular distributions and reconstruct a triangular fuzzy number with its vagueness, support, core and membership function. In other words, we were able to establish, for each professor, a relationship with all the other professors based on their subjective view of the relationship with their colleagues. This approach assumes the possibility that the relationship is asymmetric. Prof A might overestimate a collaboration with Prof. B, who, on the other hand, believes that the collaboration is of minor or significant importance. Each professor answered the question for every possible department colleague, leading us to the creation of the entire departmental network.
Figure 5 and Table 1 present the results of the fuzzy in-degree centrality indexes of the actors involved in the Department of Economic Studies of the University of Chieti-Pescara, Italy. The most central professors are 29, 44, and 21. Clearly, for privacy reasons, we do not disclose the names of the department members. The results can be evaluated both graphically and by observing the centre of gravity. Naturally, we could also consider other ranking methods among those available in the literature. Figure 6 and Table 2 represent the results of the fuzzy out-degree centrality indexes of the actors involved in the Department of Economic Studies of the University of Chieti-Pescara, Italy. The most central professors are 7, 38, and 35. It’s interesting to note how we get very different values if comparing in- and out-degree centrality indexes.
Figure 5: Fuzzy in-degree centrality indexed of the actors in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
Prof Left Core Right CoG
29 0.09 0.11 0.13 0.11
44 0.08 0.10 0.12 0.10
21 0.07 0.10 0.11 0.09
26 0.07 0.08 0.10 0.08
28 0.06 0.08 0.09 0.08
37 0.06 0.08 0.09 0.08
38 0.07 0.08 0.08 0.08
45 0.06 0.08 0.09 0.08
46 0.07 0.08 0.09 0.08
47 0.06 0.08 0.10 0.08
48 0.07 0.08 0.09 0.08
17 0.05 0.07 0.09 0.07
24 0.05 0.07 0.08 0.07
30 0.06 0.08 0.08 0.07
2 0.04 0.06 0.07 0.06
3 0.04 0.06 0.07 0.06
16 0.05 0.06 0.08 0.06
20 0.06 0.06 0.07 0.06
4 0.03 0.05 0.06 0.05
7 0.04 0.06 0.06 0.05
Table 1: First twenty Professors according to normalized Fuzzy in-degree centrality indexes in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
Figure 6: Fuzzy out-degree centrality indexed of the actors in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
Prof Left Core Right CoG
7 0.36 0.42 0.47 0.42
38 0.31 0.41 0.48 0.40
35 0.20 0.33 0.42 0.31
9 0.15 0.24 0.28 0.22
11 0.16 0.18 0.27 0.21
27 0.15 0.19 0.21 0.18
32 0.10 0.10 0.12 0.11
46 0.10 0.11 0.12 0.11
21 0.05 0.08 0.09 0.07
29 0.04 0.06 0.06 0.06
43 0.05 0.05 0.08 0.06
44 0.05 0.05 0.08 0.06
53 0.04 0.06 0.07 0.06
22 0.04 0.05 0.05 0.05
8 0.03 0.05 0.05 0.04
12 0.03 0.03 0.04 0.04
18 0.03 0.04 0.05 0.04
26 0.03 0.04 0.04 0.04
17 0.03 0.03 0.03 0.03
37 0.02 0.03 0.03 0.03
Table 2: First twenty Professors according to normalized Fuzzy out-degree centrality indexes in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
Figure 7 presents the findings of the fuzzy betweenness centrality indexes of the nodes of the Department of Economic Studies of the University of Chieti-Pescara, Italy. The most fundamental professors are 38, 35, and 7. In other words, professor 38 is often on the best path for many couple of professors because, for them, it’s better to reach other colleagues passing via professor 38.
Figure 7: Fuzzy betweenness centrality indexed of the actors in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
Figure 8, Table 3, Figure 9, Table 4, illustrate the results of the fuzzy in-closeness and out-closeness centrality indexes, respectively. The most important nodes are 7, 35, 38, and 29, 44, 21, respectively. In summary, professor 29 is the most powerful and approachable.
Figure 8: Fuzzy in-closeness centrality indexed of the actors in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
Prof Left Core Right CoG
7 0.0115 0.0131 0.0143 0.0130
35 0.0094 0.0128 0.0160 0.0127
38 0.0093 0.0131 0.0149 0.0124
22 0.0104 0.0113 0.0120 0.0112
46 0.0099 0.0110 0.0125 0.0111
32 0.0101 0.0108 0.0119 0.0109
27 0.0080 0.0110 0.0124 0.0105
11 0.0073 0.0081 0.0114 0.0089
9 0.0067 0.0087 0.0102 0.0085
17 0.0056 0.0067 0.0076 0.0066
12 0.0034 0.0042 0.0053 0.0043
21 0.0023 0.0032 0.0040 0.0032
37 0.0022 0.0024 0.0027 0.0024
8 0.0017 0.0024 0.0026 0.0022
29 0.0016 0.0023 0.0025 0.0022
44 0.0017 0.0019 0.0027 0.0021
53 0.0016 0.0021 0.0025 0.0021
49 0.0013 0.0017 0.0018 0.0016
26 0.0013 0.0015 0.0019 0.0015
10 0.0012 0.0013 0.0014 0.0013
Table 3: First twenty Professors according to normalized Fuzzy in-closeness centrality indexes in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
Figure 9: Fuzzy out-closeness centrality indexed of the actors in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
Prof Left Core Right CoG
29 0.0039 0.0045 0.0052 0.0045
44 0.0037 0.0045 0.0053 0.0045
21 0.0033 0.0039 0.0044 0.0039
28 0.0033 0.0039 0.0044 0.0039
19 0.0029 0.0035 0.0050 0.0038
41 0.0033 0.0037 0.0041 0.0037
45 0.0028 0.0035 0.0040 0.0034
23 0.0028 0.0034 0.0040 0.0034
26 0.0029 0.0033 0.0038 0.0033
36 0.0024 0.0035 0.0038 0.0033
30 0.0026 0.0033 0.0036 0.0032
48 0.0028 0.0032 0.0035 0.0032
17 0.0024 0.0030 0.0040 0.0031
47 0.0025 0.0031 0.0035 0.0031
20 0.0027 0.0028 0.0034 0.0030
16 0.0024 0.0028 0.0031 0.0028
49 0.0024 0.0028 0.0031 0.0027
8 0.0022 0.0029 0.0032 0.0027
35 0.0023 0.0027 0.0032 0.0027
39 0.0023 0.0027 0.0032 0.0027
Table 4: First twenty Professors according to normalized Fuzzy out-closeness centrality indexes in the collaborative network of the Department of Economic Studies of the University of Chieti-Pescara, Italy.
8 Discussion e conclusions
In the literature on social network analysis, the links are binary, i.e. present or absent. A possible extension of this approach, widely used in the literature, considers the so-called weighted networks. In the latter, the link between the different nodes of a network is generally expressed with a number in the range $[0,1]$. This approach of considering connections with a weight does not align with fuzzy logic’s philosophy. Fuzzy logic embraces the coexistence of opposites and, for this reason, typically conveys information about vagueness. Attempting to translate an inherently imprecise term (for example, being very good friends) into a single scalar (a number between zero and one) is a strategy that sacrifices information for simplicity.
This work considers the possibility that a vague measure can express the link between different actors in a social network. This approximate measure could be due to an attribute of human language that is often inaccurate by nature; for example, the inaccuracy inherent in concepts such as ”being friends,” ”being little friends,” ”being very close friends,” ”being close collaborators,” or ”having a poor collaboration.” In all these cases, we should have imprecise edges to preserve information about the vagueness of these links. In other cases, as in the application proposed in this work, we can also experience another kind of uncertainty due to indecision in choosing the final qualitative answer to a questionnaire, for example, by checking the mouse’s movements before providing the definitive answer.
The translation of these sources of vagueness has been addressed using triangular fuzzy numbers. This approach extends classic social networks to scenarios where actors are fully connected through nuanced ties, represented by functions. Consequently, this paper introduces a novel method for analyzing social networks, where the connections between actors are expressed through fuzzy numbers. In other words, the links between nodes are represented by functions rather than scalars, as in traditional social network theory.
In light of this new context, this study presented an extension of the traditional centrality indices to the case where the network consists of shaded edges. This research introduces several new measures that consider the imprecision of the links in calculating the centrality measures and, in particular, often provides an original interpretation of these new measures. In the application part of this work, we presented a case of a collaboration network within a university department. The goal was to measure individual collaboration through questionnaires to understand how much each pair of professors or researchers collaborates. The basic idea of this approach is to try to understand the links between department members that go beyond simple indicators such as the number of publications in common or other parameters that can be measured through scalar measures. Indeed, in a department or even in another type of network, there are many cases in which the friendship or, more generally, the collaboration between two members goes beyond the number of tangible products resulting from this collaboration. Strong cooperation between individuals can also exist simply because they have projects in common, exchange ideas, exchange didactic material, go out together outside the working environment, are relatives, and help each other with research questions even if they are not published together, have interests in common outside the university for example for professional activities, or different types of interests in common. In all these cases, considering objective parameters such as the number of joint publications would be highly limiting, and the information would need more effective collaboration between subjects. In any case, since many of these concepts we have listed are by their nature vague, the relationship between individuals and, therefore also, their centrality in a more complex network must be measured through suitable methods that consider the imprecision of human language as well as the uncertainty connected with their way of seeing things. For the latter reason, the social network we propose in this research is rarely symmetrical. In fact, in most cases, an individual’s opinion regarding his relationship with another individual is rarely the same in the opposite direction. For this reason, we have focused on the case of direct social networks in this study. Even if highly improbable, we could have as an extreme case that every link between two nodes is equal in both directions; naturally, as the size of the network or the number of nodes involved increases, the probability that this extreme case occurs tends to zero.
An exciting research component is the possibility of choosing a system of different weights to calculate the proposed centrality indices. Indeed, in some specific contexts and depending on the research objective, it could be interesting to consider other systems of weights to give different importance to solid bonds rather than weak bonds or vice versa. In the theoretical part of the research, we introduced generalized measures to allow the calculation of different centrality measures based on different possible systems of weights. In the application part, there were no particular contextual requirements to think of weights other than those all equal to each other, so we focused on a system of equal weights.
In attributing weights to calculate the centrality indices, choosing the method to create a ranking of fuzzy numbers also assumes great importance. Different methods of sorting fuzzy numbers could lead to different results. This work focused on the centre of gravity method, but our approach can be extended to other procedures. The research focused on triangular fuzzy numbers, but of course, a possible exciting extension would be to consider different membership functions to represent the links existing in a social network. Other interesting future developments could be creating an R package to implement all the functions we have made in this research and developing other indices proposed in the literature on classical social network analysis.
Our approach’s main advantage is to analyse networks of actors in which the relationships are not precise because they are defined by heterogeneous components or simply because they are collected through questionnaires whose possible answers are imprecise or simply attributes of human language. The possible applications include analysing the network of friendships on social networks such as Facebook, Instagram, YouTube, etc. Limiting oneself to simple friendships or followed contacts could be more informative. Actually, being friends with a person on Facebook can have a very nuanced meaning, depending on many other aspects that could be considered. Therefore, considering a relationship between two individuals based on a binary outcome or a simple weight between zero and one would be a complete loss of information because it does not consider the variability or, even better, the so-called fuzziness inherent in the relationship. In this research, we presented an application of the data collected within university departments, but this approach can be extended to multiple real contexts. For example, we could consider workplaces or all networks of individuals involved in sports or government teams. Consequently, the practical applications are numerous and of considerable interest.
Competing interests
No competing interest is declared.
Acknowledgments
The authors thank Dr. Simone Di Nardo for creating the data collection platform and professors Marta Di Nicola and Francesca Scozzari for supporting Dr. Annamaria Porreca in her doctoral path, during which she collected data and worked on the project. We also thank the members of the Department of Economics of the University of Pescara-Chieti for responding to the questionnaire and allowing the study to be carried out.